Calculate the Molarity (M, moles solute / Liter of solution) of a solution that is made by dissolving 3.456 µmoles of Cobalt(II) hydrogen sulfate in enough water to form a 432.1 cL solution.

Determine the Molarity (M, moles solute / Liter of solution) of a solution formed when 2.134E-3 kg Aluminum chloride is dissolved in water such that the final volume of the solution is 0.4356 kL.

Determine the Molarity (M, moles solute / Liter of solution) of a solution formed when 1.734E6 µg Barium hydroxide is dissolved in water such that the final volume of the solution is 39.56 cL.
Answer #1

1)

moles of solute = 3.456 µmoles = 3.456 x 10^-6 mol

volume = 432.1 cL = 4.321 L

Molarity = moles / volume

             = 3.456 x 10^-6 / 4.321

Molarity = 7.998 x 10^-7 M

2)

mass of AlCl3 = 2.134 x 10^-3 kg

moles of AlCl3 = 2.134 / 133.34 = 0.0160

volume = 0.4356 kL = 435.6 L

Molarity = 0.01600 / 435.6

Molarity = 3.674 x 10^-5 M